How many ounces of a 6% iodine solution needs to be added to 12 ounces of a 10% solution to create a 7% solution?
Answer by checkley77(12844)

.06x+.10*12=.07(12+x)
.06x+1.2=.84+.07x
.06x-.07x=.84-1.2
-.01x=-.36
x=-.36/-.01
x=36 ounces of 6% iodine is used.
Proof:
.06*36+.10*12=.07(12+36)
2.16+1.2=.07*48
3.36=3.36